# Efficient Adjusted Joint Significance Test and Sobel-Type

### v7

@@ 3 Multiple Mediation Testing with FWER Control
Denote $\Omega_{00}=\{k:H^{(k)}_{00}~{}{\rm holds},k=1,\cdots,d\}$, $\Omega_{10}=\{k:H^{(k)}_{10}~{}{\rm holds},k=1,\cdots,d\}$ and $\Omega_{01}=\{k:H^{(k)}_{01}~{}{\rm holds},k=1,\cdots,d\}$.
@@
The $FWER(AJS)$ is asymptotically controlled below the significance level $\delta$ for small-scale multiple testing. We emphasize that the AJS cannot be directly applied to high-dimensional mediators, such as 450K or 850K DNA methylation markers (Zhang et al., 2016; Perera et al., 2022). In the context of large-scale mediators, advanced high-dimensional statistical techniques, such as employing variable screening methods for mediator dimension reduction, are imperative, which falls beyond the scope of this paper. In this regard, we refer to the works by Zhang et al. (2016; 2021b), Huang (2019), Luo et al. (2020), Liu et al. (2022), Dai et al. (2022), Perera et al. (2022) and Tian et al. (2022), among others.
@@ 4 Adjusted Sobel-Type Confidence Intervals
$\hat{\sigma}_{\alpha\beta}^{-1}({\hat{\alpha}\hat{\beta}-\alpha\beta})\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1),$ (4.1)
where $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ denotes convergence in distribution, $\hat{\alpha}$ and $\hat{\beta}$ are the estimates for $\alpha$ and $\beta$, respectively; $\hat{\sigma}_{\alpha}$ and $\hat{\sigma}_{\beta}$ are the estimated standard errors of $\hat{\alpha}$ and $\hat{\beta}$, respectively; $\hat{\sigma}_{\alpha\beta}=\{\hat{\alpha}^{2}\hat{\sigma}^{2}_{\beta}+\hat{\beta}^{2}\hat{\sigma}^{2}_{\alpha}\}^{1/2}$. The focus of our study lies in constructing $100(1-\delta)\%$ confidence intervals for mediation effects, where $\delta$ is commonly selected as $0.05$. Based on (4.1), the Sobel-typle $100(1-\delta)\%$ confidence interval for $\alpha\beta$ is given by
@@
where $N_{1-\delta/2}(0,1)$ is the $(1-\delta/2)$-quantile of $N(0,1)$. However, the confidence interval ${\rm CI_{Sobel}}$ provided in (4.2) is excessively wide when both $\alpha$ and $\beta$ are equal to zero, resulting in a coverage probability of one for the $95\%$ confidence interval.
@@
In what follows, we provide some comparisons between ${\rm CI_{Sobel}}$ and ${\rm CI_{ASobel}}$. The Sobel-type confidence interval can be directly derived to satisfy the following expression:
@@
Hence, the asymptotic coverage probability of ${\rm CI_{ASobel}}$ is therefore closer to $(1-\delta)$ than that of ${\rm CI_{Sobel}}$ under $\alpha=\beta=0$. Furthermore, the averaged length of ${\rm CI_{ASobel}}$ in the case of $\alpha=\beta=0$, $2N_{1-\delta/2}(0,1/4)\hat{\sigma}_{\alpha\beta}$ , is significantly shorter compared to that of ${\rm CI_{Sobel}}$, which is $2N_{1-\delta/2}(0,1)\hat{\sigma}_{\alpha\beta}$. The performance of ${\rm CI_{Sobel}}$ and ${\rm CI_{ASobel}}$ will be compared through numerical simulations.
@@ 5.1 Size and Power of Single-Mediator Testing
In this section, we conduct some simulations to evaluate the performance of the AJS test for $H_{0}:\alpha\beta=0$ in the context of one mediator. Under the framework of mediation analysis, we consider three kinds of outcomes with one continuous mediator: linear mediation model (continuous outcome), logistic mediation model (binary outcome) and Cox mediation model (time-to-event outcome). To make the simulation simple and focused, we do not consider any covariates in the simulation, although our method has a similar performance when handling any covariate adjustment in these models. Specifically, the mediator $M$ is generated from the linear model $M=\alpha X+e$, where $X$ and $e$ follow from $N(0,1)$. The random outcomes are generated from the following three models:
@@
$\bullet$ Logistic mediation model$\mathbb{P}(Y=1|X,M)=\exp(\gamma X+\beta M)/\{1+\exp(\gamma X+\beta M)\}$, where $Y\in\{0,1\}$ is the binary outcome, $X$ follows from $N(0,1)$, and $\gamma=0.5$.
@@
$\bullet$ Cox mediation model: Let $T$ be the failure time, and $C$ be the censoring time. The observed survival time is $Y=\min(T,C)$. Following Cox (1972), the conditional hazard function of $T$ is $\lambda(t|X,M)=\lambda_{0}(t)\exp(\gamma X+\beta M)$, where $\lambda_{0}(t)=1$ is the baseline hazard function, $X$ and $M$ follow from $N(0,1)$; $\gamma=0.5$, $C$ is generated from $U(0,c_{0})$ with $c_{0}$ being chosen as a number such that the censoring rate is about $30\%$.
@@
Table 1: The size and power of hypothesis testing with linear mediation ${}^{\ddagger}$.
$(\alpha,\beta)$ Sobel JS ASobel AJS
$n=500$ (0, 0) 0 0.0025 0.0484 0.0487
(0, 0.5) 0.0496 0.0532 0.0496 0.0532
(0.5, 0) 0.0486 0.0526 0.0486 0.0526
(0.10, 0.15) 0.3582 0.5531 0.6984 0.7246
(0.25, 0.25) 0.9993 0.9996 0.9993 0.9996
$n=1500$ (0, 0) 0 0.0022 0.0485 0.0476
(0, 0.5) 0.0508 0.0520 0.0508 0.0520
(0.5, 0) 0.0465 0.0474 0.0465 0.0474
(0.10, 0.15) 0.9560 0.9701 0.9708 0.9777
(0.25, 0.25) 1 1 1 1
@@
From Vanderweele and Vansteelandt (2009), VanderWeele and Vansteelandt (2010) and VanderWeele (2011), the product term $\alpha\beta$ can be interpreted as the causal mediating effect of $M$ along the pathway $X\rightarrow M\rightarrow Y$ for the above three mediation models (see Figure 1). For comparison, we also use the traditional Sobel and JS methods for testing $H_{0}:\alpha\beta=0$, where the significance level is $\delta$ = 0.05. All the simulation results are based on $10000$ repetitions, where the sample size is chosen as $n=500$ and 1500, respectively. In Tables 1, S.1 and S.2, we report the sizes and powers of the Sobel, JS, ASobel and AJS methods when performing mediation tests with three kinds of outcomes. As a whole, the size of our proposed AJS and ASobel are much better than those of JS and Sobel methods, respectively. For small values of $\alpha$ and $\beta$, the power of AJS is larger than that of ASobel and the conventional Sobel and JS. These numerical findings are in line with the theoretical results of Theorem 1. In Figures 5, S.2 and S.3, we present the Q-Q plots of p-values under linear, logistic and Cox mediation models with $n=500$. The results demonstrate that the Sobel, JS, ASobel, and AJS methods accurately approximate the distribution of their respective test statistics under either $H_{10}$ or $H_{01}$. The Sobel and JS tests, however, exhibit a conservative behavior, whereas the proposed ASobel and AJS tests still accurately approximate the distribution of their corresponding test statistics. The quantiles of p-values under $H_{A}$ increase in the following order: AJS, ASobel, JS, Sobel. The aforementioned finding is consistent with the power performance presented in Table 1.
@@ 5.2 FWER and Power of Multiple-Mediators Testing
In this section, we investigate the performance of the AJS method when performing small-scale multiple testing for mediation effects via numerical simulations. The dimension of mediators is chosen as $d$ = 10, 15 and 20, respectively. Similar to Section 5.1, we consider three kinds of outcomes in the context of multiple mediators without covariates.
@@
$\bullet$ Logistic mediation model:   $\mathbb{P}(Y=1|X,\mathbf{M})=\exp(\gamma X+{\bm{\beta}}^{\prime}\mathbf{M})/\{1+\exp(\gamma X+{\bm{\beta}}^{\prime}\mathbf{M})\}$, where $Y\in\{0,1\}$ is the binary outcome, other variables and parameters are designed as those of the linear mediation model. Moreover, we consider the following settings for model parameters,
${\bm{\alpha}}$ $=$ $(0.15,0.2,0.25,0.2,0.15,0.65,0,\cdots,0)^{\prime},$
${\bm{\beta}}$ $=$ $(0.25,0.3,0.25,0.15,0.25,0,0.55,0,\cdots,0)^{\prime}.$
$\bullet$ Cox mediation model: Given the covariate $X$ and the mediator vector $\mathbf{M}$, the Cox’s conditional hazard function of failure time $T$ is $\lambda(t|X,\mathbf{M})=\lambda_{0}(t)\exp(\gamma X+{\bm{\beta}}^{\prime}\mathbf{M})$, where $\lambda_{0}(t)=1$, $\gamma=0.5$, $X$ and $\mathbf{M}$ are generated in the same way as the linear mediation model. We generate the censoring time $C$ from $U(0,c_{0})$, where $c_{0}$ is being chosen such that the censoring rate is about $30\%$. The settings of the model’s parameters are
@@
${\bm{\beta}}$ $=$ $(0.15,0.25,0.3,0.15,0.1,0,0.5,0,\cdots,0)^{\prime},$
The index set of significant mediators is $\Omega=\{1,2,3,4,5\}$. From Vanderweele and Vansteelandt (2009; 2010) and Huang and Yang (2017), the product term $\alpha_{k}\beta_{k}$ describes the causal mediation effect along the $k$th pathway $X\rightarrow M_{k}\rightarrow Y$ (see Figure 4), where $k=1,\cdots,d$. Under the significance level $\delta$, we consider the multiple testing problem $H_{0k}:~{}\alpha_{k}\beta_{k}=0$, $k=1,\cdots,d$. The proposed AJS is compared with Sobel, JS, JT_Comp Huang (2019) and DACT Liu et al. (2022) in terms of empirical FWER and Power. Note that the HDMT Dai et al. (2022) is not applicable in our simulation scenarios, as it specifically targets high-dimensional mediation hypotheses with $d\rightarrow\infty$. All the results are based on 5000 repetitions, where the sample size is $n=500$ and 1500, respectively.
In Tables 2, S.3 and S.4, we report the FWERs and Powers of five test methods in the context of linear, logistic and Cox mediation models, where the significance level is $\delta=$ 0.05. The results demonstrate that the FWER and Power of AJS outperform the other five methods significantly. The powers of all methods are observed to increase with the sample size, while they decline with the increase in dimension $d$. One explanation for this phenomenon is due to the fact that the estimated variances of model parameters are becoming larger as the increase of mediator’s dimension under fixed sample size.
@@
Dimension Sobel JS JT_Comp DACT AJS
$n=500$ $d=10$ FWER 0.0160 0.0216 0.1986 0.0006 0.0262
Power 0.0464 0.1888 0.0123 0.0242 0.3332
$d=15$ FWER 0.0108 0.0142 0.2188 0.0018 0.0308
Power 0.0303 0.1600 0.0128 0.0648 0.2964
$d=20$ FWER 0.0080 0.0108 0.2588 0.0088 0.0410
Power 0.0202 0.1402 0.0172 0.1166 0.2755
$n=1500$ $d=10$ FWER 0.0176 0.0206 0.2140 0 0.0274
Power 0.4120 0.4662 0.1241 0.0270 0.6099
$d=15$ FWER 0.0112 0.0128 0.2310 0 0.0348
Power 0.3976 0.4525 0.1518 0.2011 0.5886
$d=20$ FWER 0.0124 0.0140 0.2878 0.0030 0.0438
Power 0.3884 0.4425 0.1856 0.3392 0.5748
Table 3: The coverage probability and length of 95% confidence interval with linear mediation ${}^{\ddagger}$.
@@
$(\alpha_{k},\beta_{k})$ Sobel ASobel Sobel ASobel
$n=500$ (0, 0) 0.9998 0.9482 0.01024 0.00513
(0.35, 0) 0.9568 0.9568 0.06637 0.06637
(0.5, 0) 0.9544 0.9544 0.09450 0.09450
(0, 0.35) 0.9562 0.9562 0.06197 0.06197
(0, 0.5) 0.9530 0.9530 0.08804 0.08804
(0.15, 0.3) 0.9382 0.9380 0.06033 0.06027
(0.25, 0.35) 0.9406 0.9406 0.07723 0.07723
$n=1500$ (0, 0) 0.9998 0.9466 0.00338 0.00169
(0.35, 0) 0.9542 0.9542 0.03799 0.03799
(0.5, 0) 0.9512 0.9512 0.05410 0.05410
(0, 0.35) 0.9478 0.9478 0.03562 0.03562
(0, 0.5) 0.9536 0.9536 0.05069 0.05069
(0.15, 0.3) 0.9448 0.9448 0.03465 0.03465
(0.25, 0.35) 0.9464 0.9464 0.04415 0.04415

### v8

@@ 3 Multiple Mediation Testing with FWER Control
The $FWER(AJS)$ is asymptotically controlled below the significance level $\delta$ for small-scale multiple testing. We emphasize that the AJS cannot be directly applied to high-dimensional mediators. In the context of large-scale mediators, advanced high-dimensional statistical techniques, such as employing variable screening methods for mediator dimension reduction, are imperative, which falls beyond the scope of this paper.
@@ 4 Adjusted Sobel-Type Confidence Intervals
$\hat{\sigma}_{\alpha\beta}^{-1}({\hat{\alpha}\hat{\beta}-\alpha\beta})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}N(0,1),$ (4.1)
where $\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}$ denotes convergence in distribution, $\hat{\alpha}$ and $\hat{\beta}$ are the estimates for $\alpha$ and $\beta$, respectively; $\hat{\sigma}_{\alpha}$ and $\hat{\sigma}_{\beta}$ are the estimated standard errors of $\hat{\alpha}$ and $\hat{\beta}$, respectively; $\hat{\sigma}_{\alpha\beta}=\{\hat{\alpha}^{2}\hat{\sigma}^{2}_{\beta}+\hat{\beta}^{2}\hat{\sigma}^{2}_{\alpha}\}^{1/2}$. The focus of our study lies in constructing $100(1-\delta)\%$ confidence intervals for mediation effects, where $\delta$ is commonly selected as $0.05$. Based on (4.1), the Sobel-type $100(1-\delta)\%$ confidence interval for $\alpha\beta$ is given by
@@
where $N_{1-\delta/2}(0,1)$ is the $(1-\delta/2)$-quantile of $N(0,1)$. However, the confidence interval ${\rm CI_{Sobel}}$ provided in (4.2) is excessively wide when both $\alpha$ and $\beta$ are equal to zero. e.g., the coverage probability of the $95\%$ confidence interval provided in (4.2) approaches unity.
@@
The subsequent discussion presents a series of comparisons between ${\rm CI_{Sobel}}$ and ${\rm CI_{ASobel}}$. The Sobel-type confidence interval can be derived directly to fulfill the following expression:
@@
Hence, the asymptotic coverage probability of ${\rm CI_{ASobel}}$ is much better than that of ${\rm CI_{Sobel}}$ under $\alpha=\beta=0$. Furthermore, the averaged length of ${\rm CI_{ASobel}}$ in the case of $\alpha=\beta=0$, $2N_{1-\delta/2}(0,1/4)\hat{\sigma}_{\alpha\beta}$ , is significantly shorter compared to that of ${\rm CI_{Sobel}}$, which is $2N_{1-\delta/2}(0,1)\hat{\sigma}_{\alpha\beta}$. The performance of ${\rm CI_{Sobel}}$ and ${\rm CI_{ASobel}}$ will be compared through numerical simulations.
@@ 5.1 Size and Power of Single-Mediator Testing
In this section, we conduct some simulations to evaluate the performance of the ASobel and AJS tests for $H_{0}:\alpha\beta=0$ in the context of one mediator. Under the framework of mediation analysis, we consider three kinds of outcomes with one continuous mediator: linear mediation model (continuous outcome), logistic mediation model (binary outcome) and Cox mediation model (time-to-event outcome). Specifically, the mediator $M$ is generated from the linear model $M=\alpha X+{\bm{\eta}}^{\prime}\mathbf{Z}+e$, where $X$ and $e$ follow from $N(0,1)$, $\mathbf{Z}=(Z_{1},Z_{2})^{\prime}$ with $Z_{1}$ and $Z_{2}$ being independent random variables following $N(0,1)$, ${\bm{\eta}}=(0.5,0.5)^{\prime}$. The random outcomes are generated from the following three models:
@@
$\bullet$ Logistic mediation model: Let $Y\in\{0,1\}$ be the binary outcome, and
${P}(Y=1|X,M,\mathbf{Z})=\frac{\exp(\gamma X+\beta M+{\bm{\theta}}^{\prime}\mathbf{Z})}{1+\exp(\gamma X+\beta M+{\bm{\theta}}^{\prime}\mathbf{Z})},$
where $\gamma=0.5$ and ${\bm{\theta}}=(0.5,0.5)^{\prime}$.
@@
$\bullet$ Cox mediation model: Let $T$ be the failure time, and $C$ be the censoring time. The observed survival time is $Y=\min(T,C)$. Following Cox (1972), the conditional hazard function of $T$ is $\lambda(t|X,M,\mathbf{Z})=\lambda_{0}(t)\exp(\gamma X+\beta M+{\bm{\theta}}^{\prime}\mathbf{Z})$, where $\lambda_{0}(t)=1$ is the baseline hazard function, $\gamma=0.5$, ${\bm{\theta}}=(0.5,0.5)^{\prime}$; $C$ is generated from $U(0,c_{0})$ with $c_{0}$ being chosen as a number such that the censoring rate is about $30\%$.
@@
Table 1: The size and power of hypothesis testing with linear mediation model.
$(\alpha,\beta)$ Sobel JS Bootstrap ASobel AJS
$n=200$ (0, 0) 0 0.0016 0.0010 0.0432 0.0460
(0, 0.5) 0.0406 0.0498 0.0556 0.0406 0.0498
(0.5, 0) 0.0420 0.0522 0.0568 0.0420 0.0522
(0.15, 0.15) 0.1190 0.3068 0.2774 0.4184 0.5124
(0.25, 0.25) 0.7470 0.8814 0.7584 0.7834 0.8978
$n=500$ (0, 0) 0 0.0034 0.0008 0.0482 0.0446
(0, 0.5) 0.0540 0.0568 0.0542 0.0540 0.0568
(0.5, 0) 0.0460 0.0496 0.0546 0.0460 0.0496
(0.15, 0.15) 0.6820 0.8334 0.8218 0.8740 0.9090
(0.25, 0.25) 0.9990 0.9998 0.9946 0.9990 0.9998
$n=1000$ (0, 0) 0.0002 0.0026 0.0012 0.0498 0.0496
(0, 0.5) 0.0456 0.0482 0.0520 0.0456 0.0482
(0.5, 0) 0.0466 0.0476 0.0486 0.0466 0.0476
(0.15, 0.15) 0.9886 0.9950 0.9952 0.9954 0.9974
(0.25, 0.25) 1 1 1 1 1
Under the regularity conditions in Vanderweele and Vansteelandt (2009), VanderWeele and Vansteelandt (2010) and VanderWeele (2011), the product term $\alpha\beta$ can be interpreted as the causal mediating effect of $M$ along the pathway $X\rightarrow M\rightarrow Y$ for the three mediation models (see Figure 1). For comparison, we also use the traditional Sobel, JS and Bootstrap methods for testing $H_{0}:\alpha\beta=0$, where the significance level is $\delta$ = 0.05. The popular quantile Bootstrap test method is employed for assessing the mediation effect, and $H_{0}$ will be rejected if 0 does not fall within the interval $[Q_{\delta/2}(\{\hat{\alpha}^{(b)}\hat{\beta}^{(b)}\}_{b=1}^{B}),~{}Q_{1-\delta/2}(\{\hat{\alpha}^{(b)}\hat{\beta}^{(b)}\}_{b=1}^{B})]$, where $Q_{\delta/2}(\cdot)$ is the $(\delta/2)$-empirical quantile function, $\hat{\alpha}^{(b)}$ and $\hat{\beta}^{(b)}$ are the corresponding parameter estimators with the $b$-th Bootstrap samples, $b=1,\cdots,B$.
@@
All the simulation results are based on $5000$ repetitions, where $B=1000$ and the sample size is chosen as $n=200$, $500$ and 1000, respectively. In Tables 1, S.1 and S.2, we report the sizes and powers of the Sobel, JS, Bootstrap, ASobel and AJS methods when performing mediation tests with three kinds of outcomes. Under $\alpha=\beta=0$, the size of our proposed AJS and ASobel are much better than those of JS, Sobel and Bootstrap methods, respectively. For small values of $\alpha$ and $\beta$, the power of AJS is larger than that of ASobel and the conventional Sobel, JS and Bootstrap methods. These numerical findings are in line with the theoretical results of Theorem 1. In Figures 5, S.2 and S.3, we present the Q-Q plots of p-values under linear, logistic and Cox mediation models with $n=500$. The Bootstrap method does not depend on the use of p-values when conducting hypothesis tests for $H_{0}$. The Q-Q plots demonstrate that the Sobel, JS, ASobel, and AJS methods accurately approximate the distribution of their respective test statistics under either $H_{10}$ or $H_{01}$. The Sobel and JS tests, however, exhibit a conservative behavior, whereas the proposed ASobel and AJS tests still accurately approximate the distribution of their corresponding test statistics. The quantiles of p-values under $H_{A}$ increase in the following order: AJS, ASobel, JS, Sobel. The aforementioned finding is consistent with the power performance presented in Table 1.
@@ 5.2 FWER and Power of Multiple-Mediators Testing
In this section, we investigate the performance of the AJS method when performing small-scale multiple testing for mediation effects via numerical simulations. The dimension of mediators is chosen as $d$ = 10, 15 and 20, respectively. Similar to Section 5.1, we consider three kinds of outcomes in the context of multiple mediators.
@@
$\bullet$ Logistic mediation model:  Let $Y\in\{0,1\}$ be the binary outcome, and
${P}(Y=1|X,\mathbf{M},\mathbf{Z})=\frac{\exp(\gamma X+{\bm{\beta}}^{\prime}\mathbf{M}+{\bm{\theta}}^{\prime}\mathbf{Z})}{1+\exp(\gamma X+{\bm{\beta}}^{\prime}\mathbf{M}+{\bm{\theta}}^{\prime}\mathbf{Z})},$
where the variables $X$, $\mathbf{M}$, $\mathbf{Z}$ remain consistent with those in those of the linear mediation model, except for the settings for model parameters,
${\bm{\alpha}}$ $=$ $(0.15,0.2,0.25,0.5,0.25,0.65,0,\cdots,0)^{\prime},$
${\bm{\beta}}$ $=$ $(0.25,0.3,0.35,0.65,0.55,0,0.55,0,\cdots,0)^{\prime}.$
$\bullet$ Cox mediation model: Given the covariate $X$, the mediator vector $\mathbf{M}$ and the covariates vector $\mathbf{Z}$, the Cox’s conditional hazard function of failure time $T$ is
$\lambda(t|X,\mathbf{M},\mathbf{Z})=\lambda_{0}(t)\exp(\gamma X+{\bm{\beta}}^{\prime}\mathbf{M}+{\bm{\theta}}^{\prime}\mathbf{Z}),$
where $\lambda_{0}(t)=1$, $\gamma=0.5$, $X$, $\mathbf{M}$ and $\mathbf{Z}$ are generated in the same way as the linear mediation model. We generate the censoring time $C$ from $U(0,c_{0})$, where $c_{0}$ is being chosen such that the censoring rate is about $30\%$. The settings of the model’s parameters are
@@
${\bm{\beta}}$ $=$ $(0.15,0.25,0.3,0.15,0.1,0,0.5,0,\cdots,0)^{\prime}.$
The index set of significant mediators is $\Omega=\{1,2,3,4,5\}$. Based on Vanderweele and Vansteelandt (2009; 2010) and Huang and Yang (2017), the product term $\alpha_{k}\beta_{k}$ describes the causal mediation effect along the $k$th pathway $X\rightarrow M_{k}\rightarrow Y$ (see Figure 4), where $k=1,\cdots,d$. Under the significance level $\delta$, we consider the multiple testing problem $H_{0k}:~{}\alpha_{k}\beta_{k}=0$, $k=1,\cdots,d$. The proposed AJS is compared with Sobel, JS, JT_Comp Huang (2019), DACT Liu et al. (2022) and Bootstrap in terms of empirical FWER and Power. The Bootstrap test method will reject $H_{0k}$ if 0 does not fall within the Bonferroni corrected interval $[Q_{\delta/(2d)}(\{\hat{\alpha}_{k}^{(b)}\hat{\beta}_{k}^{(b)}\}_{b=1}^{B}),~{}Q_{1-\delta/(2d)}(\{\hat{\alpha}_{k}^{(b)}\hat{\beta}_{k}^{(b)}\}_{b=1}^{B})]$, where $\hat{\alpha}_{k}^{(b)}$ and $\hat{\beta}_{k}^{(b)}$ are the corresponding parameter estimators with the $b$-th Bootstrap samples, $b=1,\cdots,B$. The HDMT Dai et al. (2022) is not applicable in our simulation scenarios, as it specifically targets high-dimensional mediation hypotheses with $d\rightarrow\infty$. All the results are based on 5000 repetitions, where $B=1000$, the sample size is $n=200$, 500 and 1000, respectively.
In Tables 2, S.3 and S.4, we report the FWERs and Powers of six test methods in the context of linear, logistic and Cox mediation models, where the significance level is $\delta=$ 0.05. The results demonstrate that the FWER and Power of AJS outperform the other five methods significantly. The powers of all methods are observed to increase with the sample size, while they decline with the increase in dimension $d$. One explanation for this phenomenon is due to the fact that the estimated variances of model parameters are becoming larger as the increase of mediator’s dimension under fixed sample size.
@@
Dimension Sobel JS JT_Comp DACT Bootstrap AJS
$n=200$ $d=10$ FWER 0.0062 0.0192 0.1706 0.0116 0.0294 0.0232
Power 0.0011 0.0195 0.0032 0.0146 0.0127 0.0764
$d=15$ FWER 0.0052 0.0126 0.1982 0.0252 0.0176 0.0282
Power 0.0004 0.0147 0.0034 0.0236 0.0092 0.0568
$d=20$ FWER 0.0038 0.0112 0.2176 0.0486 0.0186 0.0338
Power 0.0002 0.0103 0.0042 0.0314 0.0076 0.0458
$n=500$ $d=10$ FWER 0.0166 0.0222 0.2074 0 0.0296 0.0262
Power 0.0452 0.1836 0.0099 0.0252 0.1265 0.3286
$d=15$ FWER 0.0080 0.0134 0.2282 0.0006 0.0230 0.0306
Power 0.0299 0.1560 0.0135 0.0646 0.1094 0.2954
$d=20$ FWER 0.0084 0.0108 0.2748 0.0086 0.0190 0.0388
Power 0.0216 0.1387 0.0199 0.1160 0.0965 0.2762
$n=1000$ $d=10$ FWER 0.0172 0.0198 0.2102 0 0.0268 0.0246
Power 0.3055 0.4042 0.0505 0.0279 0.3547 0.5211
$d=15$ FWER 0.0100 0.0128 0.2382 0.0002 0.0164 0.0342
Power 0.2669 0.3864 0.0605 0.1408 0.3390 0.5004
$d=20$ FWER 0.0086 0.0100 0.2748 0.0020 0.0178 0.0352
Power 0.2361 0.3734 0.0771 0.2534 0.3219 0.4857
Table 3: The coverage probability and length of 95% confidence interval with linear mediation model.
@@
$(\alpha_{k},\beta_{k})$ Sobel Bootstrap ASobel Sobel Bootstrap ASobel
$n=200$ (0, 0) 1 0.9986 0.9482 0.02665 0.04009 0.01388
(0.35, 0) 0.9676 0.9454 0.9632 0.10811 0.11234 0.10764
(0.5, 0) 0.9522 0.9460 0.9522 0.15321 0.15638 0.15321
(0, 0.35) 0.9706 0.9394 0.9614 0.10098 0.11246 0.10009
(0, 0.5) 0.9614 0.9414 0.9614 0.14158 0.15478 0.14158
(0.25, 0.35) 0.9376 0.9466 0.9376 0.12626 0.13371 0.12606
(0.45, 0.5) 0.9424 0.9450 0.9424 0.19413 0.20249 0.19413
$n=500$ (0, 0) 0.9996 0.9972 0.9548 0.01024 0.01558 0.00513
(0.35, 0) 0.9572 0.9474 0.9572 0.06678 0.06769 0.06678
(0.5, 0) 0.9520 0.9472 0.9520 0.09488 0.09527 0.09488
(0, 0.35) 0.9604 0.9498 0.9604 0.06238 0.06834 0.06237
(0, 0.5) 0.9498 0.9440 0.9498 0.08817 0.09578 0.08817
(0.25, 0.35) 0.9474 0.9484 0.9474 0.07813 0.08223 0.07813
(0.45, 0.5) 0.9424 0.9410 0.9424 0.12088 0.12599 0.12088
$n=1000$ (0, 0) 0.9996 0.9986 0.9472 0.00510 0.00776 0.00255
(0.35, 0) 0.9534 0.9454 0.9534 0.04668 0.04684 0.04668
(0.5, 0) 0.9488 0.9468 0.9488 0.06658 0.06658 0.06658
(0, 0.35) 0.9530 0.9484 0.9530 0.04376 0.04763 0.04376
(0, 0.5) 0.9480 0.9484 0.9480 0.06228 0.06755 0.06228
(0.25, 0.35) 0.9462 0.9476 0.9462 0.05483 0.05762 0.05483
(0.45, 0.5) 0.9544 0.9534 0.9544 0.08502 0.08845 0.08502